OpenStudy (anonymous):

find the vertex and axis of symmetry of the graph of the function: f(x)=x^2-8x

OpenStudy (anonymous):

help

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -(-8)/(2*1) x = 4 So the axis of symmetry is 4 Now plug this in to get f(x) = x^2 - 8x f(4) = (4)^2 - 8(4) f(4) = -16 So the vertex is (4, -16)